# A decoder-only foundation model for time-series forecasting

### v2

@@ 5 Pretraining Details
We would like our pretraining corpus to include large volumes of temporal data representing a variety of domains, trend patterns and time granularities that ideally capture the forecasting use-cases which we are interested in serving by the deployed model. It is challenging to find a large time-series dataset that meets the volume and diversity of data needed for training our foundation model. The bulk of real data used to train our models come from Google trends  [^3] https://trends.google.com and Wiki Pageview statistics  [^4] https://en.wikipedia.org/wiki/Wikipedia:Pageview_statistics. In summary the main data sources are:
@@
Google Trends. Google Trends captures search interest over time for millions of queries. We choose around 22k head queries based on their search interest over 15 years from 2007 to 2022. We use the search interest over time for these queries in hourly, daily, weekly and monthly granularities to form our dataset. The date ranges are Jan. 2018 to Dec. 2019 for hourly and Jan. 2007 to Dec. 2021 for the other granularities.
@@
Wiki Pageviews. Wiki Pageviews captures the hourly views of all Wikimedia pages. We download all pageview data from Jan. 2012 to Nov. 2023, clean and aggregate the views by page into hourly, daily, weekly and monthly granularities, and filter out the resulting time-series with excessive zeros. The final corpus contains O(100B) time points.
Other real data sources. Along with the wiki and trends data, we also add time-series from several other publicly available datasets to our pretraining corpus. We add in all the granularities of the M4 dataset [17] and the hourly (and 15 minute) Electricity and hourly Traffic datasets (see  [34]). We also add the 10 minute granularity Weather dataset used for evaluations in [34]. M4 has a good mix of granularities with around 100k time-series in total. Traffic and Electricity are large long-term forecasting datasets with $>$ 800 and $>$ 300 time-series each having tens of thousands of time-points. In addition, we add all the 15 min granularity traffic time series from [29].
@@
Synthetic. Another major component of our pretraining data is of synthetic origin. We create generators for ARMA [18] processes, seasonal patterns (mixture of sines and cosines of different frequencies), trends (linear, exponential with a few change-points) and step functions. A synthetic time-series can be an additive combination of 1 or more of these processes. We create 3M synthetic time-series each of length 2048 time-points. More details are presented in Appendix A.4.
@@
Dataset Mixing. We train on a mixture distribution over these datasets that aims to give sufficient weight to all granularities. The training mixture is 50% real data and 50% synthetic, with the real data mixture providing equal weights to all hourly + sub-hourly, daily, weekly, and monthly datasets. We train with a maximum context length of 512 whenever the length of the time-series allows that. For weekly granularity we do not have sufficiently long time-series therefore a maximum context length of 256 is used. For the same reason, a maximum context length of 64 is used while training on $\geq$ monthly granularity data.
@@
Figure 2: We report average performance in three groups of datasets. In all figures, the lower the metric the better and the error bars represent one standard error. Note that among the baselines onlye TimesFM and llmtime are zero-shot. In (a) we report results on the Monash datasets. Since the datasets have different scales, we average the MAE’s scaled by the MAE of a naive baseline. We can see that TimesFM is among the top-3 models. In (b), we report the similarly scaled MAE on the Darts benchmarks. TimesFM is within significance of the top of the best performing method which is ARIMA in this case. Note that these datasets have one time-series each and therefore statistical methods are competitive with deep learning ones. Finally, in (c) we report the average MAE for 96 and 192 horizon prediction tasks on 4 ETT datasets i.e 8 tasks in total. TimesFM and PatchTST are the best performing models in this case.
@@ 6.1 Zero-shot Evaluation
In all cases, we report performance on the official metrics and scalings of the datasets. We present a summary of the results below - more detailed experimental results can be found in Appendix A.2.
@@
Monash [8]. Monash archive is a collection of 30 datasets of different training and prediction lengths that covers granularities ranging from minutes to years and domains including finance, demand forecasting, weather and traffic. The archive reports four official metrics for several statistical baselines such as Exponential Smoothing(ETS) and ARIMA, as well as supervised ML baselines like CatBoost [23], DeepAR [25] and WaveNet [20]. Following llmtime [9] we start from the Monash Huggingface repository  [^5] https://huggingface.co/datasets/monash_tsf and filter out the datasets that contain missing values. This leaves us with 18 datasets.
@@
Out of the four official metrics, following prior work [9], we report our performance in terms of mean MAE (see Appendix A.1). As the datasets have massively different scales, for each dataset we normalize the metric by the metric achieved by a naive baseline that just constantly predicts the last value in the context for each time-series. Then the scaled MAE’s are averaged across all datasets. The scaled aggregation was also used in [9].
@@
The mean scaled MAE across all datasets is plotted in Figure 1(a) along with standard error bars. We compare the performance of TimesFMwith the baseline models implemented in Monash, and the zero-shot llmtime [9] model that uses GPT-3 [24] with a specific prompting technique. Note that the zero-shot models are marked as (Zero-Shot). TimesFM is among the top 3 models even though we never trained on these datasets. It outperforms some deep supervised models like DeepAR and FFNN, and improves on llmtime’s performance by more than 10%.
@@
Darts [10]. This is a collection of 8 univariate datasets which include interesting seasonalities and additive+multiplicative trends. We report performance of several baselines implemented in the Darts package like TCN [14], N-HiTS [4] and N-BEATS [21]. All these baselines are supervised. As before, we also report zero-shot forecasting results from llmtime [9] using GPT-3 [24]. Other supervised baselines implemented in [9] like SM-GP [31] and seasonal ARIMA [18] are also added.
@@
We report the official metric for this dataset group that is MAE for each individual dataset in Appendix A.2. In Figure 1(b), we present the average scaled MAE across all 8 datasets, as we did for the Monash datasets. TimesFM is within statistical significance of the best model that is seasonal ARIMA in this case. Note that since there are only 8 individual time-series in this dataset group, the standard errors are not sharp and therefore does not provide a clear ordering among the models.
@@
Informer [34]. The Informer datasets have been widely used for benchmarking various supervised long-horizon forecasting methods. A few of these datasets are used in pretraining, so we focus on the other datasets in this collection (ETTm1, ETTm2. ETTh1 and ETTh2) related to electricity transformer temperatures over a two year period in $1$ hour and $15$ minutes granularities. Note that the long horizon baselines usually report rolling validation results on the test set which would amount to millions of tokens for evaluating llmtime [9] and would be too expensive. Therefore, following the latter paper, we compare all methods on the last test window. Also, it is reasonable to directly average the MAE for these datasets since the results are reported on standard normalized dataset (using the statistics of the training portion).
@@
We consider the task of predicting horizon length 96 and 192, given a context length of 512 for all methods. The MAE averaged over all 8 tasks (4 datasets with two horizons each) is presented in Figure 1(b). TimesFM performs the best and the supervised PatchTST [19] baseline (which is a state-of-the-art long horizon deep forecasting method) is within significance of it. The other long horizon methods are quite a bit worse even though they have been trained these datasets. llmtime(ZS) is better than FEDFormer but worse than PatchTST in a statistically significant way.
@@ 6.2 Ablation
Scaling. Performance curves with respect to number of parameters in a model have been a keenly studied area in the context of LLMs.  Kaplan et al. [12] established a power law like relationship between the number of parameters in a language model and its downstream performance i.e the more the number of paramaters the better the performance. However,  Hoffmann et al. [11] established a more nuanced scaling law that lays down methods to train compute optimal models based on the number of tokens available in a training dataset. We perform a preliminary scaling study where we train three TimesFM models of sizes 17M, 70M and 200M parameters, using the same pretraining dataset and till the similar number of iterations.
We provide the mean scaled MAE results on the Monash datasets in Figure 3. It can be clearly seen that the errors decrease monotonically with the number of parameters, thus pointing to a similar observation as [12], but in the context of time-series.
Figure 3: Average scaled MAE on Monash datasets for three different TimesFM model sizes. We can see that the performance improves with increasing model size.
@@
Autoregressive decoding. In recent long-term forecasting works [33, 19, 6] it has been observed that directly predicting the entire forecasting horizon in one shot from a decoder can yield better results than auto-regressive decoding on long horizon benchmarks. For a foundation model the horizon length of the task is not known before inference time, therefore one-shot decoding might not be possible for very long horizons. However, by keeping the output_patch_len longer than input_patch_len one can ensure fewer autoregressive steps. This was one of the key decisions in the design of TimesFM, that is quite different from LLMs. In order to showcase this we choose the task of predicting 512 time-steps into the future for the ETT datasets. In Figure 4, we present results from a model with output_patch_len=32 vs our original model that uses output_patch_len=128. The former has to perform 16 autoregressive steps while the latter has to do only 4. It can be clearly seen that having a larger output_patch_len helps in this case.
Figure 4: Ablation with respect to output patch length for the task of predicting 512 steps into the future, in the same setting as that of Section 6.1. We report the average across all ETT 4 datasets.
@@
Input patch length.  The size of input_patch_len represents an important trade-off. We have typically seen that increasing its value from 8 to 32 increases performance but having too high a input_patch_len is impractical since that makes the model shift from decoder only training more towards encoder-decoder style training. Note that in the ”Training” paragraph of Section 4, we describe the mask sampling strategy to support any context length. If in the extreme case $p$ is set the maximum context length we have to individually sample all possible context windows from 1 to maximum context length, which would be required for encoder-decoder style of training.
In Figure 5 we show the mean scaled MAE TimesFM(ZS) - 70M model with input_patch_len=8 on Monash datasets, which is clearly worse than our original model that uses input_patch_len=32.
Figure 5: Average scaled MAE for our 70M models on Monash datasets for two different input patch lengths. We can see that the performance improves with increasing patch size.
6.3 Visualizing Forecast
We conduct a visual inspection of the forecasts generated by TimesFM(ZS) on the benchmark datasets. Our inspection in particular emphasizes time-series exhibiting both trend and seasonal components readily interpretable by humans. As illustrated in Figure 6, TimesFM(ZS) effectively captures these subtle characteristics within both the trend and seasonal patterns of the depicted time-series.
Figure 6: Forecasts visualized on Darts and Monash datasets. Plots on the right zoom in the forecasting parts from the left.
@@ 7 Conclusion
In this paper, we presented TimesFM, a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised forecasting models on a diverse set of time series data. This model is trained on real-world and synthetic datasets comprising around 100B timepoints using a patched-decoder style attention architecture with around 200M parameters.
In future work, we plan to delve into a more theoretical understanding of how such a time-series foundation model can obtain good performance for out-of-distribution data, and also investigate the fine-tuning/few-shot performance of this model.

### v4

@@ 5 Pretraining Details
We would like our pretraining corpus to include large volumes of temporal data representing a variety of domains, trend and seasonality patterns and time granularities that ideally capture the forecasting use-cases which we are interested in serving by the deployed model. It is challenging to find a large time-series dataset that meets the volume and diversity of data needed for training our foundation model. We address this problem by sourcing the bulk of data used to train our models from three major sources: Google trends, Wiki Pageview statistics and synthetic time-series. In summary the main data sources are:
@@
Google Trends. Google Trends  [^3] https://trends.google.com captures search interest over time for millions of queries. We choose around 22k head queries based on their search interest over 15 years from 2007 to 2022. Beyond these head queries the time-series become more than 50% sparse. We download the search interest over time for these queries in hourly, daily, weekly and monthly granularities to form our dataset. The date ranges are Jan. 2018 to Dec. 2019 for hourly and Jan. 2007 to Dec. 2021 for the other granularities. The trends datasets amounts to roughly 0.5B time-points.
@@
Wiki Pageviews. Wiki Pageviews  [^4] https://en.wikipedia.org/wiki/Wikipedia:Pageview_statistics captures the hourly views of all Wikimedia pages. We download all pageview data from Jan. 2012 to Nov. 2023, clean and aggregate the views by page into hourly, daily, weekly and monthly granularities, and filter out pageview time-series with excessive zeros. The final corpus contains roughly 300B time-points.
@@
Synthetic Data. Another major component of our pretraining data is of synthetic origin. We create generators for ARMA [20] processes, seasonal patterns (mixture of sines and cosines of different frequencies), trends (linear, exponential with a few change-points) and step functions. A synthetic time-series can be an additive combination of one or more of these processes. We create 3M synthetic time-series each of length 2048 time-points. More details about our synthetic data generation are presented in Appendix A.8.
Other real-world data sources. Along with the wiki and trends data, we also add time-series from several other publicly available datasets to our pretraining corpus. We add all the granularities of the M4 dataset [22], the hourly and 15 minute Electricity and the hourly Traffic datasets (see  [44]). We also add the 10-minute granularity Weather dataset used for evaluations in [44]. M4 has a good mix of granularities with around 100k time-series in total. Traffic and Electricity are large long-term forecasting datasets with $>$ 800 and $>$ 300 time-series each having tens of thousands of time-points. In addition, we add all the 15 min granularity traffic time-series from [36].
Table 1: Composition of TimesFM pretraining dataset.
Dataset Granularity # Time series # Time points
Synthetic 3,000,000 6,144,000,000
Electricity Hourly 321 8,443,584
Traffic Hourly 862 15,122,928
Weather [44] 10 Min 42 2,213,232
Favorita Sales Daily 111,840 139,179,538
LibCity [36] 15 Min 6,159 34,253,622
M4 hourly Hourly 414 353,500
M4 daily Daily 4,227 9,964,658
M4 monthly Monthly 48,000 10,382,411
M4 quarterly Quarterly 24,000 2,214,108
M4 yearly Yearly 22,739 840,644
Wiki hourly Hourly 5,608,693 239,110,787,496
Wiki daily Daily 68,448,204 115,143,501,240
Wiki weekly Weekly 66,579,850 16,414,251,948
Wiki monthly Monthly 63,151,306 3,789,760,907
Trends hourly Hourly 22,435 393,043,680
Trends daily Daily 22,435 122,921,365
Trends weekly Weekly 22,435 16,585,438
Trends monthly Monthly 22,435 3,821,760
@@
Dataset Mixing and Training. We train on a mixture distribution over these datasets that aims to give sufficient weight to all granularities and datasets. The training loader samples 80% real data and 20% synthetic, with the real data mixture providing equal weights to the groups: hourly + sub-hourly, daily, weekly, and monthly datasets. We train with a maximum context length of 512 whenever the length of the time-series allows that. For weekly granularity we do not have sufficiently long time-series; therefore a maximum context length of 256 is used. For the same reason, a maximum context length of 64 is used while training on $\geq$ monthly granularity data. We also use only the standard normalization part of reversible instance normalization [16] – i.e, the context of each time-series is scaled by the context mean and standard deviation of the first input patch in the context.
@@
Figure 2: We report average performance in three groups of datasets. In all figures, the lower the metric the better and the error bars represent one standard error. Note that among the baselines only TimesFM and llmtime are zero-shot. In (a) we report results on the Monash datasets. Since the datasets have different scales, we take the Geometric Mean (GM) of the MAE’s scaled by the MAE of a naive baseline. We can see that TimesFM is the top model. In (b), we report the similarly scaled MAE on the Darts benchmarks. TimesFM is within significance of the best performing methods which are ARIMA and llmtime in this case. Note that these datasets have one time-series each and therefore statistical methods are competitive with deep learning ones. Finally, in (c) we report the average MAE for 96 and 192 horizon prediction tasks on 4 ETT datasets i.e 8 tasks in total. TimesFM and PatchTST are the best performing models
@@ 6.1 Zero-shot Evaluation
In all cases, we report performance on the official metrics and scalings of the datasets, using either their standard test splits or common test splits in other literature. We present a summary of the results below - more details can be found in Appendix A.5. We provide the hyper-parameters and other details about our model in Appendix A.6.
@@
Monash [9]. Monash archive is a collection of 30 datasets of different training and prediction lengths that covers granularities ranging from minutes to years and domains including finance, demand forecasting, weather and traffic. The archive reports four official metrics for several statistical baselines such as Exponential Smoothing(ETS) and ARIMA, as well as supervised ML baselines like CatBoost [28], DeepAR [30] and WaveNet [27]. Following llmtime [11] we start from the Monash Huggingface repository  [^5] https://huggingface.co/datasets/monash_tsf and filter out the datasets that contain missing values. This leaves us with 18 datasets which we specify in Appendix A.5.2.
@@
Out of the four official metrics, following prior work [11], we report our performance in terms of mean MAE (see Appendix A.2). As the datasets have massively different scales, for each dataset we normalize the metric by the metric achieved by a naive baseline that just constantly predicts the last value in the context for each time-series. Then the scaled MAE’s are averaged across all datasets. The scaled aggregation was also used in [11]. In Figure 2(a), we use the Geometric Mean (GM) for averaging since it is more robust for normalized metrics [8]. We also report the Arithmetic Mean based aggregated metrics in Figure 4 in the appendix.
@@
The mean scaled MAE across all datasets is plotted in Figure 2(a) along with standard error bars. We compare the performance of TimesFM with the baseline models implemented in Monash, and the zero-shot llmtime [11] model that uses GPT-3 [29] with a specific prompting technique. Note that the zero-shot models are marked as (Zero-Shot). TimesFM is the top model even though we never trained on these datasets. It is slightly better but within significance of N-BEATS but outperforms deep supervised models like DeepAR [30], and improves on llmtime’s performance by more than 25%.
@@
Darts [14]. This is a collection of 8 univariate datasets which include interesting seasonalities and additive+multiplicative trends. We report performance of several baselines implemented in the Darts package like TCN [19], N-HiTS [5] and N-BEATS [25]. All these baselines are supervised. As before, we also report zero-shot forecasting results from llmtime [11] using GPT-3 [29]. Other supervised baselines in [11] like SM-GP [35] and ARIMA [20] are also added.
@@
We report the official metric for this dataset group that is MAE for each individual dataset in Appendix A.5. In Figure 2(b), we present the average scaled MAE across all 8 datasets, as we did for the Monash datasets. TimesFM is within statistical significance of the best models that is llmtime and seasonal ARIMA in this case. Note that since there are only 8 individual time-series in this dataset group, the standard errors are not sharp and therefore does not provide a clear ordering among the models. Also, note that for ARIMA, the seasonality needs to be encoded correctly in the parameters for the best results, which needed manual tuning. Further, since these datasets are used in numerous time series blog posts for illustrative purposes, data contamination for llmtime cannot be ruled out.
@@
Informer [44]. The Informer datasets have been widely used for benchmarking various supervised long-horizon forecasting methods. A few of these datasets are used in pretraining, so we focus on the other datasets in this collection (ETTm1, ETTm2. ETTh1 and ETTh2) related to electricity transformer temperatures over a two year period in $1$ hour and $15$ minutes granularities. Note that the long horizon baselines usually report rolling validation results on the test set which would amount to millions of tokens for evaluating llmtime [11] and would be too expensive. Therefore, following llmtime, we compare all methods on the last test window. Also, it is reasonable to directly average the MAE for these datasets since the results are reported on standard normalized dataset (using the statistics of the training portion).
@@
We consider the task of predicting horizon length 96 and 192, given a context length of 512 for all methods. The MAE averaged over all 8 tasks (4 datasets with two horizons each) is presented in Figure 2(b). TimesFM performs the best and the supervised PatchTST [24] baseline (which is a state-of-the-art long horizon deep forecasting method) is within significance of it. The other long horizon methods are quite a bit worse even though they have been trained these datasets. llmtime is better than FEDFormer but worse than PatchTST with statistical significance.
We present visual examples of our forecasts along with baselines in Appendix A.9.
@@ 6.2 Ablation
(a) Scaled MAE (GM) on Monash datasets as a function of FLOPS across three model sizes 17M, 70M and 200M. The first 4 points are from 17M and 70M checkpoints while the last 3 are from 200M.
(b) Ablation with respect to output patch length for the task of predicting 512 steps into the future on ETT datasets on the original test set in [44]. We report the average across all 4 ETT datasets.
(c) Scaled MAE (GM) for our 70M models on Monash datasets for different input patch lengths. We also plot error bars denoting one standard error.
(d) Average scaled MAE for Monash on the left and average MAE on ETT datasets on the right. We compare the performance of 200M model with and without the synthetic data.
Figure 3: Ablation studies with respect to various design choices.
@@
Scaling. Performance curves with respect to number of parameters in a model have been a keenly studied area in the context of LLMs.  [17] established a power law like relationship between the number of parameters in a language model and its downstream performance i.e the more the number of paramaters the better the performance. However,  [13] established a more nuanced scaling law that lays down methods to train compute optimal models based on the number of tokens available in a training dataset.
We perform a preliminary scaling study where we train three TimesFM models of sizes 17M, 70M and 200M parameters, using the same pre-training dataset till 1.5M iterations with a global batch-size of 4096. Then we collect checkpoints that represent varying number of FLOPS (Floating Point OPerationS) across the different model runs. Then we plot the performance on Scaled MAE (GM) on Monash as a function of FLOPS, in Figure 3(a). This is now a standard way to perform scaling studies in LLMs (see recent work like [10]). It can be clearly seen that the errors decrease monotonically with the number of FLOPS (in log scale). All experiments were performed on a TPUv5e [^6] https://cloud.google.com/tpu/docs/v5e-training setup with 16 tensor-cores. For the 200M model it takes 2 days to complete 1.5M iterations on our setup.
@@
Autoregressive Decoding. In recent long-term forecasting works [40, 24, 7] it has been observed that directly predicting the entire forecasting horizon in one shot from a decoder can yield better results than auto-regressive decoding on long horizon benchmarks. For a foundation model, the horizon length of the task is not known before inference time, therefore one-shot decoding might not be possible for very long horizons. However, as mentioned earlier, by keeping the output_patch_len longer than input_patch_len one can ensure fewer autoregressive steps. This was one of the key decisions in the design of TimesFM, that is quite different from LLMs. In order to showcase this we choose the task of predicting 512 time-steps into the future for the ETT datasets on the original rolling validation task of the ETT test sets [44]. In Figure 3(b), we present results from models with output_patch_len varying from 8 to 128. We see a monotonic decrease in average MAE with output_patch_len.
@@
Input Patch Length.  The size of input_patch_len represents an important trade-off. We have typically seen that increasing its value from 8 to 32 increases performance but having too high a input_patch_len is impractical since that makes the model shift from decoder only training more towards encoder-decoder style training. Note that in the "Training" paragraph of Section 4, we describe the mask sampling strategy to support any context length. If in the extreme case $p$ is set the maximum context length we have to individually sample all possible context windows from 1 to maximum context length, which would be required for encoder-decoder style of training.
In Figure 3(c), we show the mean scaled MAE (GM) TimesFM(ZS) - 70M model on Monash with input_patch_len varying from 8 to 128. Note that both models have been trained to about 1.5M steps even though the p=8 model is three times slower to train. We can see that $p=16,32$ marks the best performance, with the error increasing towards either end. Note that $p=32$ model is almost twice as fast to train compared to $p=16$ and thus constitutes a prudent choice.
Dataset Ablation. Next we showcase the need for synthetic data. Intuitively, the majority of our real datasets have commonly found granularities like hourly, daily etc which have specific periodic patterns like 24 time-point period for hourly data. This can make the model not generalize well to underrepresented frequencies. We train a 200M model with no synthetic data added in the mix and showcase the performance on Monash and ETT datasets in Figure 3. It can be seen that there is a performance drop on Monash because many of the datasets in Monash have under-represented granularities like quarterly, yearly or 10 minutes etc. Perhaps even more compelling is the comparison on ETT datasets. We can see that there is almost no difference between the two models on the hourly ETTh datasets which has a well represented granularity. However, for the 15min ETTm datasets the model with synthetic data performs quite a bit better.
We provide a finetuning study in the same setting as [42] in Appendix A.3, where our model performs better than all baselines on all the reported datasets. This shows the utility of our model on downstream tasks.
@@ 7 Conclusion
In this paper, we presented TimesFM, a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised forecasting models on a diverse set of time-series data. This model is pretrained on real-world and synthetic datasets comprising O(100B) timepoints. We discuss limitations and future work in more detail in Appendix A.1.
8 Impact Statement
This paper shows that it is possible to train a single pretrained model that has phenomenal zero-shot performance on a variety of forecasting tasks, thus opening up exciting possibilities for downstream applications. Therefore it is crucial to discuss ethical and societal considerations of using such a model and how some of the related concerns can be mitigated.
Data Privacy. Note that most of our data sources are publicly available and are aggregated i.e no individual user activity constitutes a time-point. Further the Google Trends data is differentially private.
Bias. Biases can creep into a foundation model through a variety of sources especially through data. The model might perpetuate these biases in forecasts, leading to unfair outcomes. Biased forecasts can have real-world consequences. For example, a biased forecast of crime rates in a neighborhood could lead to increased police presence, disproportionately impacting certain communities. Note that our model is not trained with any covariates so some of these sensitivities are reduced but cannot be ruled out.
In this regard, we believe that it is best to release the exact details of the datasets used for training and therefore we have summarized our data sources in Table 1. Moreover we plan to do a open weights release of our model so that the community can analyze the model for downstream tasks. We will ensure that we release a good model card [23]. This will also aid in finetuning the model with more diverse data-sources.
Training cost. Our largest model has 200M parameters which is much smaller compared to even smaller scale LLMs. Our data volume is quite large but still smaller compared to SOTA LLMs. We have revealed the exact computation requirements for the models i.e 16 core TPUv5e for 2 days. Note that experimentation and trial runs of course cost more than the final model run. Therefore in the interest of equitability we would like to release the weights of our model in a responsible manner.
Lastly we would like to note that similar to LLMs there could be inputs on which the model does not perform well or hallucinates. Therefore in many critical use cases it might be recommended to use the model in a human-in-the-loop fashion or alternatively do a wide range of testing or finetuning.